K-10. If a and b are integers, 30, 310,5 ta, and 56, prove that a = (mod 15).
Answer #1

3 ta → a= 3 kt l l = 1,2 then a² = 3 x 7 e ² + 6kl but l = 1,2 l = 1, 4 so a² = 1 (mod 3) a = 1 (mod 3) for any such a simila